OpenStudy (anonymous):

Ð1 and Ð2 are congruent. If mÐ1 = 10x – 5 and mÐ2 = 6x + 15, then what is the degree measure of Ð1? Select one: a. 45° b. 20° c. 30° d. 25°

OpenStudy (johnweldon1993):

If they are congruent they will equal each other, so we can just set these equal to each other and solve for 'x' \[\large D1 = D2\] \[\large 10x - 5 = 6x + 15\] \[\large 4x = 20\] \[\large x = 5\] Now if that is true...since \[\large D1 = 10x - 5\] Puggin n the 5 for 'x' we get that \[\large D1 = 10(5) - 5 = 50 - 5 = 45\]

OpenStudy (anonymous):

so its a ?

OpenStudy (johnweldon1993):

Correct :)

OpenStudy (anonymous):

thanks . i appreciate it (: do you think you could help me with a couple more ?